# Enhancing PLS of Indoor IRS-VLC Systems for Colluding and Non-Colluding Eavesdroppers

### v1

@@ Abstract
Most intelligent reflecting surface (IRS)-aided indoor visible light communication (VLC) studies ignore the time delays introduced by reflected paths, even though these delays are inherent in practical wideband systems. In this work, we adopt a realistic assumption of IRS-induced time delay for physical layer security (PLS) enhancement. We consider an indoor VLC system where an IRS is used to shape the channel so that the reflected signals add constructively at the legitimate user and create intersymbol interference at eavesdroppers located inside the coverage area. The resulting secrecy capacity maximisation over the IRS element allocation is formulated as a complex combinatorial optimisation problem and is solved using deep reinforcement learning with proximal policy optimisation (PPO). The approach is evaluated for both colluding eavesdroppers, which combine their received signals, and non-colluding eavesdroppers, which act independently. Simulation results are shown for various simulation setups, which demonstrate significant secrecy capacity gains. In a worst-case scenario, where the eavesdroppers have stronger channels than the legitimate user, the proposed PPO-based IRS allocation improves secrecy capacity by 107% and 235% in the colluding and non-colluding cases, respectively, compared with allocating all IRS elements to the legitimate user. These results demonstrate that time-delay-based IRS control can provide a strong secrecy advantage in practical indoor VLC scenarios.
I Introduction
@@
vlc has attracted significant interest as a key enabling technology for future sixth generation (6G) wireless networks. The rapidly growing demand for high speed wireless connectivity and the increasing congestion and licensing cost of the radio frequency (RF) spectrum motivate the search for alternative access technologies, and visible light communication (VLC) addresses this demand by exploiting the large optical band between approximately 400 THz and 790 THz [haas2016lifi]. In addition to this abundant bandwidth, VLC offers several attractive features, including licence free operation, high energy efficiency through light emitting diode (LED) luminaires, low electromagnetic interference, and the potential for seamless integration with existing solid state lighting infrastructure [elgala2011indoor, karunatilaka2015ledvlcsurvey]. These properties make VLC particularly well suited to high density indoor environments such as offices, hospitals, and industrial facilities, where reliable, secure, and fast throughput connectivity is required.
@@ II-B NLoS Channel
III Secrecy Performance Analysis
@@
III-A Secrecy Capacity for Non-Colluding Eavesdroppers
@@
III-B Secrecy Capacity for Colluding Eavesdroppers
@@
IV Optimisation and Learning Framework for IRS-Aided Secrecy Enhancement
@@
IV-A Secrecy Capacity Maximisation Problem
@@
IV-B Proposed PPO Based Solution
@@
TABLE I: PPO Hyperparameters
Parameter Value
Training episodes $N_{\text{ep}}$ 1500
Batch size $N_{\text{batch}}$ 64
Clipping parameter $\epsilon$ 0.2
GAE parameter $\lambda$ 0.95
Discount factor $\gamma$ 0.99
Actor learning rate $\alpha_{\pi}$ $1\times 10^{-4}$
Critic learning rate $\alpha_{V}$ $5\times 10^{-4}$
Hidden layer dimension 256
V Simulation Results
@@
TABLE II: Room layout, user positions, and IRS configuration
Parameter Value
@@
Number of IRS elements $N_{\text{irs}}$ $15\times 15$
Transmit optical power $P_{\text{opt}}$ 11 W to $10$ W
Bob position (best case) $(2.5,2.5,0.75)$ m
Eve1 position (best case) $(4.5,4.5,0.75)$ m
Eve2 position (best case) $(4.0,4.0,0.75)$ m
Bob position (worst case) $(4.5,4.5,0.75)$ m
Eve1 position (worst case) $(2.5,2.5,0.75)$ m
Eve2 position (worst case) $(2.0,2.0,0.75)$ m
TABLE III: System parameters
Parameter Value
LED half power semi angle $\Phi_{\scriptscriptstyle 1/2}=60^{\circ}$
PD responsivity [A/W] $R_{\scriptscriptstyle\text{pd}}=0.6$
PD physical area [cm2] $A_{\scriptscriptstyle d}=1$
PD FoV $\Psi_{\scriptscriptstyle\text{FoV}}=90^{\circ}$
Symbol period [ns] $T_{s}=1$
Noise power spectral density $\mathcal{N}=10^{-21}$
IRS reflectivity $\rho=1$
Refractive index of PD lens $\xi=1.5$
Modulation scaling factor $\mathsf{k}=3.2$
Gap value [dB] $\Gamma=2$
Element separation [cm] $D=30$

### v2

@@ Abstract
Most intelligent reflecting surface (IRS)-aided indoor visible light communication (VLC) studies ignore the time delays introduced by reflected paths, even though these delays are inherent in practical wideband systems. In this work, we adopt a realistic assumption of IRS-induced time delay for physical layer security (PLS) enhancement. We consider an indoor VLC system where an IRS is used to shape the channel so that the reflected signals add constructively at the legitimate user and create intersymbol interference at eavesdroppers located inside the coverage area. The resulting secrecy capacity maximisation over the IRS element allocation is formulated as a complex combinatorial optimisation problem and is solved using deep reinforcement learning with proximal policy optimisation (PPO). The approach is evaluated for both colluding eavesdroppers, which combine their received signals, and non-colluding eavesdroppers, which act independently. Simulation results are presented for a range of indoor layouts and user placements, demonstrating significant secrecy capacity gains. In a worst-case scenario, where the eavesdroppers have stronger channels than the legitimate user, the proposed PPO-based IRS allocation improves secrecy capacity by 107% and 235% in the colluding and non-colluding cases, respectively, compared with allocating all IRS elements to the legitimate user. These results demonstrate that time-delay-based IRS control can provide a strong secrecy advantage in practical indoor VLC scenarios.
{IEEEkeywords}
Visible light communication, intelligent reflecting surface, physical layer security, time delay, wideband channels, secrecy capacity, colluding eavesdroppers, non-colluding eavesdroppers, proximal policy optimisation.
1 INTRODUCTION
@@
Vsible light communication (VLC) has attracted significant interest as a key enabling technology for future sixth generation (6G) wireless networks. The rapidly growing demand for high speed wireless connectivity and the increasing congestion and licensing cost of the radio frequency (RF) spectrum motivate the search for alternative access technologies, and VLC addresses this demand by exploiting the large optical band between approximately 400 THz and 790 THz [haas2016lifi]. In addition to this abundant bandwidth, VLC offers several attractive features, including licence free operation, high energy efficiency through light emitting diode (LED) luminaires, low electromagnetic interference, and the potential for seamless integration with existing solid state lighting infrastructure [elgala2011indoor, karunatilaka2015ledvlcsurvey]. These properties make VLC particularly well suited to high density indoor environments such as offices, hospitals, and industrial facilities, where reliable, secure, and fast throughput connectivity is required.
@@ 2.2 NLoS Channel
3 Secrecy Performance Analysis
@@
3.1 Secrecy Capacity for Non-Colluding Eavesdroppers
@@
3.2 Secrecy Capacity for Colluding Eavesdroppers
@@
4 Optimisation and Learning Framework for IRS-Aided Secrecy Enhancement
@@
4.1 Secrecy Capacity Maximisation Problem
@@
4.2 Proposed PPO Based Solution
@@
Table 1: PPO Hyperparameters
Symbol Quantity Value
$N_{\text{ep}}$ Training episodes 1500
$N_{\text{batch}}$ Batch size 64
$\epsilon$ Clipping parameter 0.2
$\lambda$ GAE parameter 0.95
$\gamma$ Discount factor 0.99
$\alpha_{\pi}$ Actor learning rate $1\times 10^{-4}$
$\alpha_{V}$ Critic learning rate $5\times 10^{-4}$
$d_{h}$ Hidden layer dimension 256
5 Simulation Results
@@
Table 2: Room layout, user positions, and IRS configuration
Symbol Quantity Value
@@
$N_{\text{irs}}$ Number of IRS elements $15\times 15$
$P_{\text{opt}}$ Transmit optical power 11 W to $10$ W
Best case
$\mathbf{p}_{\text{B}}$ Bob position $(2.5,2.5,0.75)$ m
$\mathbf{p}_{\text{E}_{1}}$ Eve1 position $(4.5,4.5,0.75)$ m
$\mathbf{p}_{\text{E}_{2}}$ Eve2 position $(4.0,4.0,0.75)$ m
Worst case
$\mathbf{p}_{\text{B}}$ Bob position $(4.5,4.5,0.75)$ m
$\mathbf{p}_{\text{E}_{1}}$ Eve1 position $(2.5,2.5,0.75)$ m
$\mathbf{p}_{\text{E}_{2}}$ Eve2 position $(2.0,2.0,0.75)$ m
Table 3: System parameters
Symbol Quantity Value
$\Phi_{\scriptscriptstyle 1/2}$ LED half-power semi-angle $60^{\circ}$
$R_{\scriptscriptstyle\text{pd}}$ PD responsivity [A/W] $0.6$
$A_{\scriptscriptstyle d}$ PD physical area [cm2] 11
$\Psi_{\scriptscriptstyle\text{FoV}}$ PD FoV $90^{\circ}$
$T_{s}$ Symbol period [ns] 11
$\mathcal{N}$ Noise power spectral density $10^{-21}$
$\rho$ IRS reflectivity 11
$\xi$ Refractive index of PD lens $1.5$
$\mathsf{k}$ Modulation scaling factor $3.2$
$\Gamma$ Gap value [dB] 22
$D$ Element separation [cm] $30$
@@ 6 Conclusion
{IEEEbiography}
[[Uncaptioned image]] Rashid Iqbal (Graduate Student Member, IEEE) received his B.Sc. degree in Electrical Engineering from CECOS University of IT and Emerging Sciences, Peshawar, Pakistan, in 2019. He earned his M.Sc. degree in Electrical and Electronic Engineering from the Institute of Space Technology (IST), Islamabad, Pakistan, in 2022. He is pursuing a Ph.D. in Electrical and Electronic Engineering at the University of Glasgow, UK (2023–2026). His research interests include indoor visible light communication (VLC), physical layer security (PLS), optical intelligent reflecting surfaces (OIRS), integrated sensing and communication (ISAC), channel modeling, and optimization techniques for wireless communications.
{IEEEbiography}
[[Uncaptioned image]] Ahmed Zoha (Senior Member, IEEE) received the PhD degree from the esteemed 6G/5GIC Centre, University of Surrey, Guildford, U.K., in 2014. He is currently an associate professor with the James Watt School of Engineering, University of Glasgow, Glasgow, U.K. He has more than 15 years of experience. His research interests include AI, machine learning, and advanced signal processing. He was the recipient of three prestigious IEEE best paper awards. He was endorsed as an exceptional U.K. talent by the Royal Academy of Engineering and awarded to early-career world-leading innovators and scientists.
{IEEEbiography}
[[Uncaptioned image]] Salama Ikki (Senior Member, IEEE) is currently a Professor and the Research Chair of wireless communications with Lakehead University, Thunder Bay, ON, Canada. He is the author of more than 100 journals and conference papers and has more than 8000 citations and an H-index of 41. His research group has made substantial contributions to 4G and 5G wireless technologies. His group’s current research interests include massive MIMO, cell-free massive MIMO, visible light communications, and wireless sensor networks. He was a recipient of several awards for research, teaching, and services.
{IEEEbiography}
[[Uncaptioned image]] MOHAMMAD Ali IMRAN (Fellow, IEEE) is a Professor of Communication Systems, Dean of Transnational Engineering Education, Dean of Graduate Studies for College of Science and Engineering and head of the Communications, Sensing and Imaging hub at the University of Glasgow, UK. He also holds affiliate professor positions with the University of Oklahoma, USA, and the 5G/6G Innovation Centre of the University of Surrey, UK. He has a grant portfolio of £20M+, published 600 journal and conference papers, and edited/published 13 books. He is a Fellow of IEEE, the Royal Society of Arts, the Royal Society of Edinburgh, the European Alliance of Innovation, and the Institution of Engineering and Technology.
{IEEEbiography}
[[Uncaptioned image]] Dr Hanaa Abumarshoud (Senior Member, IEEE) is a Lecturer (Assistant Professor) in the James Watt School of Engineering at the University of Glasgow, UK. Prior to this, she was a Postdoctoral Research Associate at the LiFi Research and Development Centre at the University of Strathclyde (2020–2022) and the University of Edinburgh (2017–2020). She received her MSc and PhD in Electrical and Computer Engineering from Khalifa University, UAE, in 2013 and 2017, respectively. Dr Abumarshoud was endorsed as a “Global Talent” by the Royal Academy of Engineering in the UK in 2022 and was named among the “100 Brilliant and Inspiring Women in 6G” by the Women in 6G community in both 2024 and 2025 for her contributions to the field. She is a Senior Member of IEEE and served as an Associate Editor for IEEE Communications Letters (2022-2024). She regularly contributes as an Organising Committee Member for flagship conferences, organises and chairs technical workshops, and serves as a reviewer for leading international journals and conferences within IEEE, the Royal Society, and the Optical Society. She received the IEEE TAOS Best Paper Award in 2022.